OpenStudy (anonymous):

what is a equivalent fraction for 120 over 132?

OpenStudy (anonymous):

120 / 132 = 60 / 66 = 30/33 = 10/11

OpenStudy (anonymous):

120 / 132 = 12 * 10 / 12 * 11 = 10/11